# Task-Oriented Communications for Visual Navigation with Edge-Aerial Collaboration in Low Altitude Economy

### v1

@@ IV Performance Evaluation
Fig. 7 depicts end-to-end latency (encoding, transmission, and decoding) of each compression scheme under three communication bottleneck constraints (4, 8 and 12 KB/s). At 4 KB/s, O-VIB with fixed feature coordinates $d=32$ and $d=128$ complete in 5.1 ms and 25.1 ms, respectively, whereas WebP, JPEG, H.265 and H.264 incur 5.8 s, 9.1 s, 7.2 s and 10.0 s, respectively. As throughput increases, O-VIB latency decreases roughly in proportion to reduced transmission time, while all codecs remain above one second. Moreover, enlarging the latent dimension $d$ raises neural encoding/decoding overhead and payload size, thereby increasing total delay. These findings confirm that embedding orthogonality in VIB yields ultra-compact, task-oriented features and achieves over three orders-of-magnitude lower latency under stringent bandwidth constraints.

### v3

@@ IV Performance Evaluation
Fig. 7 shows end-to-end latency (encoding, transmission, decoding) under 4, 8, and 12 KB/s bottlenecks (poor channel conditions). At 4 KB/s, O-VIB achieves $0.24$ s ($d=32$) and $0.85$ s ($d=128$), while WebP, H.265, H.264, and JPEG incur $5.7$ s, $7.1$ s, $10.9$ s, and $9.1$ s. Compared to WebP, O-VIB reduces latency by $95.7\%$. At 8 KB/s, O-VIB drops to $0.15$ s ($d=32$) and $0.44$ s ($d=128$), while WebP remains at $2.9$ s, achieving a $95.0\%$ reduction. At 12 KB/s, O-VIB further lowers latency to $114.0$ ms and $0.31$ s, compared to WebP’s $1.9$ s, realizing a $94.2\%$ reduction. These results confirm that O-VIB maintains sub-second latency and achieves over an order of magnitude improvement under stringent bottlenecks.